OpenStudy (anonymous):

You pull with a force of 300 N on a piece of luggage of mass 49 kg, but it does not move. If the coefficient of friction between the luggage and the floor is 0.95, what is the force of static friction acting on the luggage? A. 500 N B. 456 N C. 300 N D. 656 N

OpenStudy (shamim):

Given m=49kg Coefficient of friction=0.95

OpenStudy (shamim):

Weight of49 kg=mg=49*9.8=?

OpenStudy (shamim):

Static frictional force= coefficient of static friction*R=0.95*49*9.8=?

OpenStudy (shamim):

Because R=weight=49*9.8

OpenStudy (shamim):

U know R is a force acting on the object by the floor

OpenStudy (shamim):

Weight is a force acting on the floor by luggage

OpenStudy (shamim):

U know Every action has an equal n oposit reaction

OpenStudy (shamim):

So weight has also an equal n oposit reaction

OpenStudy (shamim):

So Weight=R

OpenStudy (shamim):

Weight n R oposit too

OpenStudy (shamim):

But we take magnitude only

OpenStudy (shamim):

Response plz!!!!!!!

OpenStudy (anonymous):

Ok got it!

OpenStudy (shamim):

Glsd to help u!!!!!

OpenStudy (shamim):

*glad